A hot lump of 30.9 g of copper at an initial temperature of 97.4 °C is placed in 50.0 mL of H2O initially at 25.0 °C and allowed to reach thermal equilibrium. What is the final temperature of the copper and water given that the specific heat of copper is 0.385 J/(g·°C)? Assume no heat is lost to surroundings.
Answer #1

qmetal   = qwater

mCmetal\DeltaT = mCwater\DeltaT

30.9 x (0.385) (97.4 -X) = 50 x (4.18) (X - 25)

11.896 (97.4 - X) = 209 (X - 25)

1158.67 - 11.896X = 209X - 5225

220.896 X = 6383.67

X = 28.9

final temperature = 28.9 0C